OpenStudy (anonymous):

A corn refining company produces corn gluten cattle feed at a variable cost of $82 per ton. If fixed costs are $120 000 per month and the feed sells for $134 per ton, how many tonnes must be sold each month for the company to have a monthly profit of $560 000?

OpenStudy (amistre64):

C(x) = 82x - 120000 R(x) = 134x P(x) = R(x) - C(x); then derive and maximize

OpenStudy (amistre64):

.. its not a max problem :) 560000 = R(x) - C(x)